Question 11 Find the derivative: f(x) = x2 In 5x 2x (3x In 5x) X+ In 10x **Previous
Answer #1

We have the product rule in differentiation as follows.

If f = u.v then f' = u.v' + u'.v .

Now we have f(x) = x^{2} ln 5x .

Thus by the product rule we get

\begin{align*} f'(x) & = x^{2} (ln 5x)' + (x^{2})' ln 5x \\ & = x^{2}. (\frac{1}{5x} 5) + 2x. ln 5x \\ & = \frac{x^{2}}{x} + 2x. ln 5x \\ & = x + 2x. ln 5x\\ &=x(1+2ln5x) \end{align*}.